OpenStudy (anonymous):

Your rocket takes off from Earth and goes from 0 m per second to 560 m per second in the first 7 seconds of flight. What is your acceleration?

OpenStudy (anonymous):

a=(v-u)/t a=(560-0)/7 a=80m/s^2

OpenStudy (anonymous):

can u help me on my other acceleration problems i have 3 others?

OpenStudy (anonymous):

go on..

OpenStudy (anonymous):

Once your rocket reaches space, the captain adjusts the speed. Your captain goes from 24,000 km/hr to 17,000 km/hr over 90 seconds. What is your acceleration? Once you land on Mars, you get in your rover to begin exploring the surface. Your rover accelerates at a pace of 3.5 km/hr/second for 10 seconds. What is the velocity of your rover after 10 seconds? Light travels at 300,000 km per second. Mars is 56 million km away from Earth. How many seconds does it take light to travel from Mars to Earth?